OpenStudy (anonymous):

Suppose X and Y have joint density f(x,y)=1 for 0

OpenStudy (anonymous):

I'm confused on how to set up the limits for the double integral in this problem

OpenStudy (anonymous):

So you're looking for the probability X*Y<=z, not X<=z and Y<=z, right?

OpenStudy (anonymous):

Yes that's correct

OpenStudy (anonymous):

I'll take a look at this in a minute.

OpenStudy (anonymous):

Ok, so you're integrating over a square.

OpenStudy (anonymous):

Yes exactly, a square with corners, (0,0) (0,1) (1,1) (1,0) The distribution function F(x,y) = xy for 0≤x≤1, and 0≤y≤1 I just am confused where z comes into play for the double integral.

OpenStudy (anonymous):

The idea is that for the outer integral you integrate over the entire region, so here from 0 to 1. The variable over which you integrate in the outer integral is fixed in the inner integral. So for the inner integral you integrate over those values for which the condition holds, here XY<=Z

OpenStudy (anonymous):

Let's integrate over y in the outer integral, so we get: \[\int_0^1\int_a^b1dxdy\]

OpenStudy (anonymous):

Now to find a and b.

OpenStudy (anonymous):

Still with me so far?

OpenStudy (anonymous):

Yes, I got that far, but I was thinking a = 0, and b = z/y? But this gives a ln(0) which is why I'm confused

OpenStudy (anonymous):

That's what I had in mind too, I see the problem.

OpenStudy (anonymous):

The problem is that z/y can become bigger than 1.

OpenStudy (anonymous):

so you have to use min{1,z/y}?

OpenStudy (anonymous):

That does work, but integrating that might be hard.

OpenStudy (anonymous):

Yeah I wouldn't know how to integrate that

OpenStudy (anonymous):

I might have an idea how to integrate that: z/y<=1 y>=z So you integrate 1 from 0 to z and integrate z/y from z to 1.

OpenStudy (anonymous):

\[\int_0^1\text{min}(1,\frac{z}{y})dy=\int_0^z1dy+\int_z^1\frac{z}{y}dy\]

OpenStudy (anonymous):

Is it that simple by just adding them together like that for the integral of a min{} ?

OpenStudy (anonymous):

\[\int_0^1\text{min}(1,\frac{z}{y})dy=\int_0^z\text{min}(1,\frac{z}{y})dy+\int_z^1\text{min}(1,\frac{z}{y})dy\]

OpenStudy (anonymous):

That's true more clearly, and on the interval [0,z] 1 is smaller than z/y, and on the interval [z,1] z/y is the smaller one.

OpenStudy (anonymous):

Final answer is z-zln(z), which behaves properly: http://www.wolframalpha.com/input/?i=Plot [x-xln%28x%29%2C{x%2C0%2C1}]

OpenStudy (anonymous):

Ok Thank you. I will work through this right now and see if I get the same conclusion

OpenStudy (anonymous):

Perfect, Thanks for all your help

OpenStudy (anonymous):

You're welcome.